The pH 10 buffer used involves mixing ammonia and ammonium chloride. What is the molar ratio...

  1. The pH 10 buffer used involves mixing ammonia and ammonium chloride. What is the molar ratio of ammonia to ammonium chloride needed to produce a buffer with a pH of 10?
Answer #1

PH = 10

POH = 14-PH

          = 14-10

           = 4

PKb = 4.75

POH   = PKb + log[NH4Cl]/[NH3]

4       = 4.75 + log[NH4Cl]/[NH3]

log[NH4Cl]/[NH3] = 4-4.75  

log[NH4Cl]/[NH3]    = -0.75

[NH4Cl]/[NH3]       = 10^-0.75

[NH4Cl]/[NH3]         = 0.17783

[NH3]/[NH4Cl]     = 1/0.17783

[NH3]/[NH4Cl]     = 5.6233 /1

the molar ratio of ammonia to ammonium chloride = 5.6233 : 1